# Refining Concentration for Gaussian Quadratic Chaos

### v2

@@ III Proof of Proposition 1
The indispensable starting point in proving HW inequality is the exponential Markov inequality which gives
@@
Since $\mathbb{E}[e^{s\lambda_{i}\boldsymbol{\xi}^{2}_{i}}]=\frac{1}{\sqrt{1-2s\lambda_{i}}}$ for $12s\lambda_{i}<$ 1, we get
@@
for every $s\geq 0$ such that $12s\lambda_{i}<$ 1 for all $1\leq i\leq n$. Let $a,b>0$ be such that (7) holds. The bound in (99) can be loosened as
@@
Let $m\geq 1$ be an integer. For every $10<b<$ 1, the smallest $a$ (tightest upper bound) for which (7) holds is $a=\theta_{m}(b)$ as defined in (8). Equivalently, for every $a>\frac{1}{m+1}$, the largest $b$ for which the inequality in (7) holds is $\theta_{m}^{-1}(a)$ where $\theta_{m}^{-1}(\cdot)$ is the inverse function to $\theta_{m}(\cdot)$. Conversely, if $a\leq\frac{1}{m+1}$, there exists no $b>0$ such that (7) holds.
@@
for arbitrary $10<b<$ 1. Since $\frac{1}{8\theta_{1}(b)}$ is decreasing in $b$ and $\frac{b}{4}$ is increasing in $b$, the largest value for their minimum is achieved when $\frac{1}{8\theta_{1}(b)}=\frac{b}{4}$, i.e., $12b\theta_{1}(b)=$ 1. This equation is solved for $b\approx 0.583$ and the achieved maximum value is approximately $0.1457$ as shown in Fig. 10.
@@ V Proof of Proposition 3
To see how (14) implies HW inequality, let us write HW inequality in (2) as
@@
Interestingly, $-\frac{\Lambda(t,a)}{\min\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\}}$ depends on $t,\alpha,\beta$ only through the ratio $\rho=\frac{\alpha t}{\beta}$ in (36). For convenience in comparing the right sides in (17) and (127), we fix $a$ such that $c=1$, i.e., $\frac{3a-2}{2(1-a)^{2}}=1$. This quadratic equation has only one root in the admissible interval $\frac{2}{3}<a<1$. We denote this value of $a$ by $a_{0}$ given by
@@
Finally, we address LM inequality itself. To see why (4) and (5) are equivalent, call $t^{\prime}=2\sqrt{\beta t}+2\alpha t$. Solving for $t$, we get $t=(\frac{\sqrt{\beta+2\alpha t^{\prime}}-\sqrt{\beta}}{2\alpha})^{2}$. Relabeling $t^{\prime}$ as $t$, LM inequality in (4) becomes
@@
This inequality implies HW inequality in (124) with
@@
VI Proof of Proposition 4
We begin with the bound in (99). By Lemma 1 in Section III, for every integer $m\geq 1$ and $10<b<$ 1, the inequality in (20) is the tightest for $a=\theta_{m}(b)$. Then
@@ VI Proof of Proposition 4
for every $s\geq 0$ such that $|2s\lambda_{i}|\leq b$ for all $1\leq i\leq n$, or equivalently, $0\leq s\leq\frac{b}{2\|A\|}$. The value of $s$ that minimizes the right side of (139) is not tractable for arbitrary integer $m$. Instead, let $s_{0}$ be the value of $s$ such that the sum of the first two terms in the exponent, i.e., $-st+a2^{m}s^{m+1}\|A\|_{m+1}^{m+1}$ is minimized. It is given by
@@
$\sum_{k=2}^{m}\frac{2^{k-1}\hat{s}_{opt}^{k}\mathrm{Tr}(A^{k})}{k}\leq\sum_{k=2}^{m}\frac{2^{k-1}(\frac{b}{2\|A\|})^{k}|\mathrm{Tr}(A^{k})|}{k}=\sum_{k=2}^{m}\frac{b^{k}}{2k}\frac{|\mathrm{Tr}(A^{k})|}{\|A\|^{k}}\leq\frac{n}{2}\sum_{k=2}^{m}\frac{b^{k}}{k},$ (142)
@@
Next, we bound the sum of the first two terms in the exponent, i.e., $-\hat{s}_{opt}t+a2^{m}\hat{s}_{opt}^{m+1}\|A\|_{m+1}^{m+1}$. We consider two cases.
@@
The proof of Proposition 4 is now complete.
VII Proof of Proposition 5
@@
[…]
$\alpha,\beta$ and $\gamma$ are defined in (25) and $a=\theta_{2}(b)$ for $10<b<$ 1. The function $f(s)$ achieves it global minimum value over $s\geq 0$ at $s=s_{0}$ given by
@@
    $-ux+vx^{2}+wx^{3}\leq\frac{1}{27w^{\frac{1}{2}}w^{\prime\frac{3}{2}}}(\sqrt{v^{2}+3uw^{\prime}}-v)(v\sqrt{v^{2}+3uw^{\prime}}-v^{2}-6uw^{\prime}).$ (153)
@@
    Regarding the third term on the right side of (156), note that $s_{0}\leq\frac{\sqrt{\beta^{2}}+\sqrt{12a\gamma t}-\beta}{12a\gamma}=\sqrt{\frac{t}{12a\gamma}}$ due to the simple inequality $\sqrt{x+y}\leq\sqrt{x}+\sqrt{y}$ for $x,y\geq 0$. Then
@@
where the last step is due to the fact that $\eta_{1}(t),\eta_{2}(t)\geq 0$. Finally, we select $10<b<$ 1 such that $\min\big{\{}\frac{\sqrt{3}}{12(\theta_{2}(b))^{\frac{1}{2}}},\frac{b}{3}\big{\}}$ is maximized. This happens when $\frac{\sqrt{3}}{12(\theta_{2}(b))^{\frac{1}{2}}}=\frac{b}{3}$ or equivalently, $\frac{4}{\sqrt{3}}b(\theta_{2}(b))^{\frac{1}{2}}=1$ as promised.
@@ VIII Proof of Proposition 6
The $m_{\infty}$-bound is tighter than the relaxed HW inequality if and only if
@@
    Denoting $x=\frac{r}{n}$, we show that $\ln(1+x)<(1-2\kappa)x$ for every $x>1$. Define the function $f(x)=(1-2\kappa)x-\ln(1+x)$. Then $f^{\prime}(x)=1-2\kappa-\frac{1}{1+x}>1-2\kappa-\frac{1}{2}=\frac{\sqrt{17}-1}{16}>0$ where we used $-\frac{1}{1+x}>-\frac{1}{2}$ for $x>1$. Since $f(1)=1-2\kappa-\ln 2\approx 0.002>0$ and $f$ is increasing over $[1,\infty)$, it follows that $f(x)>0$ for every $x>1$ as desired.
Next, we look at the relaxed LM inequality in (30). Writing the bound in terms of $r=\frac{t}{\|A\|}$ and expanding the complete square in the exponent, we have
@@
Therefore, the upper bound in the $m_{\infty}$-inequality in (163) is less than the upper bound in the relaxed LM inequality if and only if
@@
But, this is true for all $r>0$ thanks to the inequality $\ln(1+x)<\sqrt{1+2x}-1$ for every $x>0$. To see this, let $f(x)=\sqrt{1+2x}-1-\ln(1+x)$. Then $f^{\prime}(x)=\frac{1}{\sqrt{1+2x}}-\frac{1}{1+x}$. But, $\sqrt{1+2x}<1+x$ for every $x>0$. As such, $f^{\prime}(x)>0$ for every $x>0$. Since $f(0)=0$ and $f$ is increasing over $[0,\infty)$, then $f(x)>f(0)=0$ for every $x>0$.
@@ IX Proof of Proposition 7
Since all coefficients are positive in $p(x+50)$, we conclude that $p(x+50)\geq 6051808>0$ for all $x\geq 0$ and hence, $p(x)>0$ for all $x\geq 50$ as desired. [^19] One can check that $p(49)=-303155<0$. This is why we addressed $4\leq m\leq 49$ and $m\geq 50$ separately.
Next, we prove that $r_{n},r^{\prime}_{n}$ are the only two positive roots for the polynomial $P_{n}^{(1)}(r)-P_{n}^{(2)}(r)$. Let us recall Descartes’ rule of signs. It says that for a polynomial $a_{0}+a_{1}x+a_{2}x^{2}+\cdots+a_{d}x^{d}$ of degree $d$, the number $N$ of positive roots and the number $N^{\prime}$ of sign changes in the coefficients have the same parity and $N\leq N^{\prime}$. The number $N^{\prime}$ is the number of indices $10\leq j\leq d-$ 1 such that $a_{j}a_{j+1}<0$. The rest of proof is devoted to showing that there are only two sign changes in the coefficients of $P_{n}^{(1)}(r)-P_{n}^{(2)}(r)$ for every even integer $n\geq 8$. Since we already showed the existence of at least two positive roots $r_{n},r^{\prime}_{n}$, it follows by Descartes’ rule of signs that $r_{n},r^{\prime}_{n}$ are the only two positive roots for $P_{n}^{(1)}(r)-P_{n}^{(2)}(r)$. Let us first determine the coefficients of $P_{n}^{(1)}(r)$ and $P_{n}^{(2)}(r)$. In case of $P_{n}^{(1)}(r)$, they are given by
@@
($\frac{{\frac{n}{2}\choose j}}{n^{j}}<\frac{1}{\sqrt{e}}\sum_{i=j}^{\frac{n}{2}-1}\frac{{i\choose j}}{2^{i}i!}$ (199)
$\stackrel{{\scriptstyle(*)}}{{\iff}}$ ($\frac{{\frac{n}{2}\choose j}}{n^{j}}<\frac{1}{\sqrt{e}}\sum_{i=0}^{\frac{n}{2}-j-1}\frac{{i+j\choose j}}{2^{i+j}(i+j)!}$
@@
where in ()(*)( ∗ ) we have changed the index $i$ to $i+j$. Call $m=\frac{n}{2}$. Then we need to show
[…]
@@
    =$\stackrel{{\scriptstyle(b)}}{{=}}$ $1-\frac{\frac{1}{2}}{m-j}$
@@ X Concluding Remarks
We made progress in improving HW inequality by increasing the absolute constant $\kappa$ in its formula from the best currently known value of $0.125$ to at least $0.145$ in the general symmetric case. This was achieved by determining the best quadratic upper bound on the function $-\ln(1-x)$ around $x=0$. In the positive-semidefinite case, it was shown that LM inequality implies HW inequality with $\kappa=1-\frac{\sqrt{3}}{2}$. We verified that indeed a larger value $\kappa=\frac{9-\sqrt{17}}{32}$ is possible through improving LM inequality. This improvement came about in two forms called the augmented LM inequality and the optimal LM inequality. The latter requires computing the unique root of a quintic polynomial in the interval $(\frac{2}{3},1)$ which must be done numerically. The former approximates this unique root by a closed-form expression where the underlying approximation error was demonstrated to be less than $0.035$ regardless of the positive-semidefinite matrix $A$ and the tail parameter $t$.
@@
We explored beyond HW inequality for general symmetric matrix $A$ by identifying the best polynomial upper bound of arbitrary degree $m+1$ on the function $-\ln(1-x)$ around $x=0$ where $m\geq 1$. As a result, a sequence of concentration bounds indexed by $m$ was developed where the case $m=1$ recovers HW inequality. These bounds are written in terms of Schatten norms of $A$ and it was conjectured that they undergo a phase transition in the sense that for given $A$, there exists a threshold $t_{c}$ such that if $t<t_{c}$, then $m=1$ (HW inequality) is the tightest and if $t>t_{c}$, then letting $m$ grow to infinity results in the sharpest bound. This led to what we called the $m_{\infty}$-bound which admits a very simple formula and it depends on $A$ only through the dimension $n$ and its operator norm $\|A\|$. An estimate (upper bound) was also derived on $t_{c}$.
@@
As mentioned earlier, every upper bound of index $m\geq 2$ in that derived sequence is looser than HW inequality if $t$ is sufficiently small. We raised the question whether one can improve upon HW inequality for all sufficiently small tail parameter $t$. This question was answered positively by finding a concentration bound that depends on all norms $\|A\|,\|A\|_{2}$ and $\|A\|_{3}$. It was called a twin to HW inequality due to the fact that its exponent scales similarly to the exponent in HW inequality, however, with a larger coefficient such that it outperforms HW inequality for both small and large values of $t$.
@@
Finally, we considered a scenario where the positive-semidefinite matrix $A$ is known only through the dimension $n$ and its largest eigenvalue (operator norm). We presented four upper bounds on the tail probability of the Gaussian quadratic chaos, namely, the $m_{\infty}$-bound, relaxed versions of HW and LM bounds and what we called the $\chi^{2}$-bound. All of these bounds depend on $A$ and $t$ through the ratio $\frac{t}{\|A\|}$. It was proved that the $m_{\infty}$-bound is sharper than the relaxed HW and LM bounds for every $n\geq 1$, $A$ and $t>0$. If $n=2,4,6$, then the $\chi^{2}$-bound is tighter than the $m_{\infty}$-bound. If $n$ is an even integer greater than or equal to $8$, then there exist positive numbers $r_{n},r^{\prime}_{n}$ such that the $m_{\infty}$-bound is tighter than the $\chi^{2}$-bound if and only if $r_{n}<\frac{t}{\|A\|}<r^{\prime}_{n}$. Moreover, as the dimension $n$ grows, $r_{n}$ moves towards zero and $r^{\prime}_{n}$ grows to infinity. It was also shown that the so-called large deviations upper bound is too loose to be useful in the sense that it is weaker than the $m_{\infty}$-bound for every $n\geq 2$, $A$ and $t>0$.
@@
    The only first-order critical numbers for $f$ are 00 and $x_{0}=1-\frac{1}{(m+1)a}$. If $a\leq\frac{1}{m+1}$, then one easily checks that $f^{\prime}(x)<0$ over the open interval $(0,1)$. Since $f(0)=0$, the Mean Value Theorem (MVT) implies that $f(x)<0$ over $(0,1)$ and hence, (7) can not hold for any $b>0$. Conversely, let $a>\frac{1}{m+1}$. Then $10<x_{0}<$ 1, $f^{\prime}(x)>0$ over the interval $(0,x_{0})$ and $f^{\prime}(x)<0$ over the interval $(x_{0},1)$. Note that $\lim_{x\to 1^{-}}f(x)=-\infty$. This tells us that $f$ rises above 00 on right of $x=0$, reaches a maximum value at the critical number $x_{0}$ and then goes down from there on and escapes to $\infty$- ∞ on left of $x=1$. As such, $f$ has a unique zero (x-crossing) somewhere over the interval $(x_{0},1)$. We let $b$ be this zero. Then (7) clearly holds for $0\leq x\leq b$. Finally, we show that $a=\theta_{m}(b)$. We have $f(b)=0$, i.e.,
    $-\ln(1-b)=b+\frac{b^{2}}{2}+\frac{b^{3}}{3}+\cdots+\frac{b^{m}}{m}+ab^{m+1}.$ (209)
    Recalling the Maclaurin series for $-\ln(1-x)$, we have $-\ln(1-b)=\sum_{i=1}^{\infty}\frac{b^{i}}{i}$. Then (209) is simplified as
    $\sum_{i=m+1}^{\infty}\frac{b^{i}}{i}=ab^{m+1}$ (210)
    and hence,
    $a=\sum_{i=m+1}^{\infty}\frac{b^{i-m-1}}{i}.$ (211)
    Changing the index $i$ to $i+m+1$, we get $a=\sum_{i=0}^{\infty}\frac{b^{i}}{i+m+1}$, i.e., $a=\theta_{m}(b)$ as promised.
@@
    Since $m$ is even, once again we have $f^{\prime}(x)<0$ for every $x<0$ and MVT enters once again to confirm that $f(x)>0$ for $x<0$.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
=$\stackrel{{\scriptstyle(b)}}{{=}}$ $e^{\frac{\epsilon}{2}}\Big{(}1-\frac{\frac{1+\epsilon}{2}+(\frac{1+\epsilon}{2})^{2}}{m^{2}}\Big{)}$

### v3

@@ III Proof of Proposition 1
The indispensable starting point in proving HWI is the exponential Markov inequality which gives
@@
Since $\mathbb{E}[e^{s\lambda_{i}\boldsymbol{\xi}^{2}_{i}}]=\frac{1}{\sqrt{1-2s\lambda_{i}}}$ for $2s\lambda_{i}<1$, we get
@@
for every $s\geq 0$ such that $2s\lambda_{i}<1$ for all $1\leq i\leq n$. Let $a>0$ and $0<b<1$ be such that (10) holds. The bound in (98) can be loosened as
@@
Let $m\geq 1$ be an integer. For every $0<b<1$, the smallest $a$ (tightest upper bound) for which (23) holds is $a=\theta_{m}(b)$ as defined in (11). Equivalently, for every $a>\frac{1}{m+1}$, the largest $b$ for which the inequality in (10) holds is $\theta_{m}^{-1}(a)$ where $\theta_{m}^{-1}(\cdot)$ is the inverse function to $\theta_{m}(\cdot)$. Conversely, if $a\leq\frac{1}{m+1}$, there exists no $b>0$ such that (10) holds.
@@
for arbitrary $0<b<1$. Since $\frac{1}{8\theta_{1}(b)}$ is decreasing in $b$ and $\frac{b}{4}$ is increasing in $b$, the largest value for their minimum is achieved when $\frac{1}{8\theta_{1}(b)}=\frac{b}{4}$, i.e., $2b\theta_{1}(b)=1$. This equation is solved for $b\approx 0.583$ and the achieved maximum value is approximately $0.1457$ as shown in Fig. 9.
@@ V Proof of Proposition 3
To see how (17) implies HWI, let us write HWI in (2) as
@@
Interestingly, $\frac{\Lambda(t,a)}{\min\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\}}$ depends on $t,\alpha,\beta$ only through the ratio $\rho=\frac{\alpha t}{\beta}$ in (40). For convenience in comparing the right sides in (20) and (126), we fix $a$ such that $c=1$, i.e., $\frac{3a-2}{2(1-a)^{2}}=1$. This quadratic equation has only one root in the admissible interval $\frac{2}{3}<a<1$. We denote this value of $a$ by $a_{0}$ given by
@@
Finally, we address LMI itself. To see why (7) and (8) are equivalent, call $t^{\prime}=2\sqrt{\beta t}+2\alpha t$. Solving for $t$, we get $t=(\frac{\sqrt{\beta+2\alpha t^{\prime}}-\sqrt{\beta}}{2\alpha})^{2}$. Relabeling $t^{\prime}$ as $t$, LMI in (7) becomes
@@
This inequality implies HWI in (123) with
@@
VI Proof of Theorem 1
We begin with the bound in (98). By Lemma 1 in Section III, for every integer $m\geq 1$ and $0<b<1$, the inequality in (23) is the tightest for $a=\theta_{m}(b)$. Then
@@ VI Proof of Theorem 1
for every $s\geq 0$ such that $|2s\lambda_{i}|\leq b$ for all $1\leq i\leq n$, or equivalently, $0\leq s\leq\frac{b}{2\|A\|}$. The value of $s$ that minimizes the right side of (138) is not tractable for arbitrary integer $m$. Instead, let $s_{0}$ be the value of $s$ such that the sum of the first two terms in the exponent, i.e., $st+a2^{m}s^{m+1}\|A\|_{m+1}^{m+1}$ is minimized. It is given by
@@
$\sum_{k=2}^{m}\frac{2^{k-1}\hat{s}_{opt}^{k}\mathrm{Tr}(A^{k})}{k}$ $\leq$ $\sum_{k=2}^{m}\frac{2^{k-1}(\frac{b}{2\|A\|})^{k}|\mathrm{Tr}(A^{k})|}{k}$ (141)
$=$ $\sum_{k=2}^{m}\frac{b^{k}}{2k}\frac{|\mathrm{Tr}(A^{k})|}{\|A\|^{k}}$
$\leq$ $\frac{n}{2}\sum_{k=2}^{m}\frac{b^{k}}{k},$
@@
Next, we bound the sum of the first two terms in the exponent, i.e., $\hat{s}_{opt}t+a2^{m}\hat{s}_{opt}^{m+1}\|A\|_{m+1}^{m+1}$. We consider two cases.
@@
The proof of Theorem 1 is now complete.
VII Proof of Theorem 2
@@
$f(s)=-st+\beta s^{2}+4a\gamma s^{3},\,\,\,\,\,0\leq s\leq\frac{b}{2\alpha},$ (149)
$\alpha,\beta$ and $\gamma$ are defined in (28) and $a=\theta_{2}(b)$ for $0<b<1$. The function $f(s)$ achieves its global minimum value over $s\geq 0$ at $s=s_{0}$ given by
@@
    $\hskip 28.45274pt-ux+vx^{2}+wx^{3}\leq\frac{1}{27w^{\frac{1}{2}}w^{\prime\frac{3}{2}}}(\sqrt{v^{2}+3uw^{\prime}}-v)(v\sqrt{v^{2}+3uw^{\prime}}-v^{2}-6uw^{\prime}).$ (152)
@@
    Regarding the third term on the right side of (155), note that $s_{0}\leq\frac{\sqrt{\beta^{2}}+\sqrt{12a\gamma t}-\beta}{12a\gamma}=\sqrt{\frac{t}{12a\gamma}}$ due to the trivial inequality $\sqrt{x+y}\leq\sqrt{x}+\sqrt{y}$ for $x,y\geq 0$. Then
@@
where the last step is due to the fact that $\eta_{1}(t),\eta_{2}(t)\geq 0$. Finally, we select $0<b<1$ such that $\min\big\{\frac{\sqrt{3}}{12(\theta_{2}(b))^{\frac{1}{2}}},\frac{b}{3}\big\}$ is maximized. This happens when $\frac{\sqrt{3}}{12(\theta_{2}(b))^{\frac{1}{2}}}=\frac{b}{3}$ or equivalently, $\frac{4}{\sqrt{3}}b(\theta_{2}(b))^{\frac{1}{2}}=1$ as promised.
@@ VIII Proof of Proposition 4
The $m_{\infty}$-bound is tighter than the relaxed HWI if and only if
@@
    Denoting $x=\frac{r}{n}$, we show that $\ln(1+x)<(1-2\kappa)x$ for every $x>1$. Define the function $f(x)=(1-2\kappa)x-\ln(1+x)$. Then $f^{\prime}(x)=1-2\kappa-\frac{1}{1+x}>1-2\kappa-\frac{1}{2}=\frac{\sqrt{17}-1}{16}>0$ where we used $\frac{1}{1+x}>-\frac{1}{2}$ for $x>1$. Since $f(1)=1-2\kappa-\ln 2\approx 0.002>0$ and $f$ is increasing over $1,\infty)$, it follows that $f(x)>0$ for every $x>1$ as desired.
Next, we look at the relaxed LMI in (33). Writing the bound in terms of $r=\frac{t}{\|A\|}$ and expanding the complete square in the exponent, we have
@@
Therefore, the upper bound in the $m_{\infty}$-inequality in (162) is less than the upper bound in the relaxed LMI if and only if
@@
But, this is true for all $r>0$ thanks to the inequality $\ln(1+x)<\sqrt{1+2x}-1$ for every $x>0$. To see this, let $f(x)=\sqrt{1+2x}-1-\ln(1+x)$. Then $f^{\prime}(x)=\frac{1}{\sqrt{1+2x}}-\frac{1}{1+x}$. But, $\sqrt{1+2x}<1+x$ for every $x>0$. As such, $f^{\prime}(x)>0$ for every $x>0$. Since $f(0)=0$ and $f$ is increasing over $0,\infty)$, then $f(x)>f(0)=0$ for every $x>0$.
Finally, we verify that the $m_{\infty}$-bound always beats the large deviations bound. Call $x=\frac{1+r}{n}\geq 1$. The bound in (94) is looser than that in (89) if and only if
$\iff$ $\Big(1+\frac{nx-1}{n}\Big)^{\frac{n}{2}}<\frac{1}{\sqrt{e}}(ex)^{\frac{n}{2}}$
$\stackrel{{\scriptstyle(*)}}{{\iff}}$ $1+\frac{nx-1}{n}<\frac{1}{e^{\frac{1}{n}}}ex$
$\iff$ $1-\frac{1}{n}+x<e^{1-\frac{1}{n}}x$
$\iff$ $1-\frac{1}{n}<(e^{1-\frac{1}{n}}-1)x,$
where in ()(*) we raised both sides to the power of $\frac{2}{n}$. Using the inequality $z<e^{z}-1$ for $z\neq 0$, we have $1-\frac{1}{n}<e^{1-\frac{1}{n}}-1$ for every $n\geq 2$. Therefore, (170) is certainly true due to $x\geq 1$.
@@ IX Proof of Proposition 5
Since all coefficients are positive in $p(x+50)$, we conclude that $p(x+50)\geq 6051808>0$ for all $x\geq 0$ and hence, $p(x)>0$ for all $x\geq 50$ as desired. [^19] One can check that p(49)=303155<0p(49)=-303155<0. This is why we addressed $4\leq m\leq 49$ and $m\geq 50$ separately.
Next, we prove that $r_{n},r^{\prime}_{n}$ are the only two positive roots for the polynomial $P_{n}^{(1)}(r)-P_{n}^{(2)}(r)$. Let us recall Descartes’ rule of signs. It says that for a polynomial $a_{0}+a_{1}x+a_{2}x^{2}+\cdots+a_{d}x^{d}$ of degree $d$, the number $N$ of positive roots and the number $N^{\prime}$ of sign changes in the coefficients have the same parity and $N\leq N^{\prime}$. The number $N^{\prime}$ is the number of indices $0\leq j\leq d-1$ such that $a_{j}a_{j+1}<0$. The rest of proof is devoted to showing that there are only two sign changes in the coefficients of $P_{n}^{(1)}(r)-P_{n}^{(2)}(r)$ for every even integer $n\geq 8$. Since we already showed the existence of at least two positive roots $r_{n},r^{\prime}_{n}$, it follows by Descartes’ rule of signs that $r_{n},r^{\prime}_{n}$ are the only two positive roots for $P_{n}^{(1)}(r)-P_{n}^{(2)}(r)$. Let us first determine the coefficients of $P_{n}^{(1)}(r)$ and $P_{n}^{(2)}(r)$. In case of $P_{n}^{(1)}(r)$, they are given by
@@
$\frac{{\frac{n}{2}\choose j}}{n^{j}}<\frac{1}{\sqrt{e}}\sum_{i=j}^{\frac{n}{2}-1}\frac{{i\choose j}}{2^{i}i!}$ (199)
$\stackrel{{\scriptstyle(*)}}{{\iff}}$ $\frac{{\frac{n}{2}\choose j}}{n^{j}}<\frac{1}{\sqrt{e}}\sum_{i=0}^{\frac{n}{2}-j-1}\frac{{i+j\choose j}}{2^{i+j}(i+j)!}$
@@
where in ()(*) we have changed the index $i$ to $i+j$. Call $m=\frac{n}{2}$. Then we need to show
$\frac{(m-1)(m-2)\cdots(m-j+1)}{m^{j-1}}<\frac{1}{\sqrt{e}}\sum_{i=0}^{m-j-1}\frac{1}{2^{i}i!},\,\,\,\,\,2\leq j\leq m-1,$ (200)
@@
    $\stackrel{{\scriptstyle(b)}}{{=}}$ $1-\frac{\frac{1}{2}}{m-j}$
@@ X Concluding Remarks
We made progress in improving HWI by increasing the absolute constant $\kappa$ in its formula from the best currently known value of $0.125$ to at least $0.145$ in the general symmetric case. This was achieved by determining the best quadratic upper bound on the function $\ln(1-x)$ around $x=0$. In the positive-semidefinite case, it was shown that LMI implies HWI with $\kappa=1-\frac{\sqrt{3}}{2}$. We verified that indeed a larger value $\kappa=\frac{9-\sqrt{17}}{32}$ is possible through improving LMI. This improvement came about in two forms called the augmented LMI and the optimal LMI. The latter requires computing the unique root of a quintic polynomial in the interval $(\frac{2}{3},1)$ which must be done numerically. The former approximates this unique root by a closed-form expression where the underlying approximation error was demonstrated to be less than $0.035$ regardless of the positive-semidefinite matrix $A$ and the tail parameter $t$.
@@
We explored beyond HWI for general symmetric matrix $A$ by identifying the best polynomial upper bound of arbitrary degree $m+1$ on the function $\ln(1-x)$ around $x=0$ where $m\geq 1$. As a result, a sequence of concentration bounds indexed by $m$ was developed where the case $m=1$ recovers HWI. These bounds are written in terms of Schatten norms of $A$ and it was conjectured that they undergo a phase transition in the sense that for given $A$, there exists a threshold $\tau_{c}$ such that if $t<\tau_{c}$, then $m=1$ (HWI) is the tightest and if $t>\tau_{c}$, then letting $m$ grow to infinity results in the sharpest bound. This led to what we called the $m_{\infty}$-bound which admits a very simple formula and it depends on $A$ only through the dimension $n$ and its operator norm $\|A\|$. An estimate (upper bound) was also derived on $\tau_{c}$.
@@
As mentioned earlier, every upper bound of index $m\geq 2$ in that derived sequence is looser than HWI if $t$ is sufficiently small. We raised the question whether one can improve upon HWI for all sufficiently small tail parameter $t$. This question was answered positively by finding a concentration bound that depends on all norms $\|A\|,\|A\|_{2}$ and $\|A\|_{3}$. It was called a twin to HWI due to the fact that its exponent scales similarly to the exponent in HWI, however, with a larger coefficient such that it outperforms HWI for both small and large values of $t$.
@@
Finally, we considered a scenario where the positive-semidefinite matrix $A$ is known only through the dimension $n$ and its largest eigenvalue (operator norm). We presented five upper bounds on the tail probability of the Gaussian quadratic chaos, namely, the $m_{\infty}$-bound, relaxed versions of HW and LM bounds, the $\chi^{2}$-bound and the large deviations bound. All of these bounds depend on $A$ and $t$ through the ratio $\frac{t}{\|A\|}$. It was proved that the $m_{\infty}$-bound is sharper than the relaxed HW and LM bounds for every $n\geq 1$, $A$ and $t>0$. The $m_{\infty}$-bound is also sharper than the large deviations bound for every $n\geq 2$, $A$ and $t>0$. If $n=2,4,6$, then the $\chi^{2}$-bound is tighter than the $m_{\infty}$-bound. If $n$ is an even integer greater than or equal to $8$, there exist positive numbers $r_{n},r^{\prime}_{n}$ such that the $m_{\infty}$-bound is tighter than the $\chi^{2}$-bound if and only if $r_{n}<\frac{t}{\|A\|}<r^{\prime}_{n}$. Moreover, as the dimension $n$ grows, $r_{n}$ moves towards zero and $r^{\prime}_{n}$ grows to infinity. It was also shown that the so-called large deviations upper bound is too loose to be useful in the sense that it is weaker than the $m_{\infty}$-bound for every $n\geq 2$, $A$ and $t>0$.
@@
    The only first-order critical numbers for $f$ are 0 and $x_{0}=1-\frac{1}{(m+1)a}$. If $a\leq\frac{1}{m+1}$, then one easily checks that $f^{\prime}(x)<0$ over the open interval $(0,1)$. Since $f(0)=0$, the Mean Value Theorem (MVT) implies that $f(x)<0$ over $(0,1)$ and hence, (10) can not hold for any $b>0$. Conversely, let $a>\frac{1}{m+1}$. Then $0<x_{0}<1$, $f^{\prime}(x)>0$ over the interval $(0,x_{0})$ and $f^{\prime}(x)<0$ over the interval $(x_{0},1)$. Note that $\lim_{x\to 1^{-}}f(x)=$-$\infty$. This tells us that $f$ rises above 0 on right of $x=0$, reaches a maximum value at the critical number $x_{0}$ and then goes down from there on and escapes to -$\infty$ on left of $x=1$. As such, $f$ has a unique zero (x-crossing) somewhere over the interval $(x_{0},1)$. We let $b$ be this zero. Then (23) clearly holds for $0\leq x\leq b$. Moreover, the equation $f(b)=0$ gives $a=\theta_{m}(b)$ as defined in (11).
@@
    Since $m$ is even, we have $f^{\prime}(x)<0$ for every $x<0$ and MVT applies once again to confirm that $f(x)>0$ for $x<0$.
@@
$\hskip 28.45274ptf(x)=x+\frac{x^{2}}{2}+\frac{x^{3}}{3}+\cdots+\frac{x^{m}}{m}+\frac{x^{m+1}}{(m+1)(1-ax)}+\ln(1-x),\,\,\,\,\,0\leq x<1.$ (210)
@@
$\forall_{\infty}m\geq 4,\,\,\,2^{m}<\frac{1}{\sqrt{e}}\frac{(1+2m)^{m-1}}{2^{m-1}(m-1)!}=\frac{1}{\sqrt{e}}\frac{(m+\frac{1}{2})^{m-1}}{(m-1)!}.$ (228)
@@
$\forall_{\infty}m\geq 4,\,\,\,2^{m}<\frac{1}{\sqrt{e}}\frac{(m+\frac{1}{2})^{m-1}e^{m-2}}{(m-1)^{m-\frac{1}{2}}}.$ (231)
@@
$\forall_{\infty}m\geq 4,\,\,\,2^{m}<\frac{1}{\sqrt{e}}\frac{e^{m-2}}{(m+\frac{1}{2})^{\frac{1}{2}}}.$ (232)
@@
$\stackrel{{\scriptstyle(b)}}{{=}}$ $e^{\frac{\epsilon}{2}}\Big(1-\frac{\frac{1+\epsilon}{2}+(\frac{1+\epsilon}{2})^{2}}{m^{2}}\Big)$